OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5)

OpenStudy (anonymous):

do you know the point slope formula?

OpenStudy (anonymous):

y-y1 = m( - x1) sooo slope (or m) = change in y divided by change in x 5--4/2--1 = 9/3 = 3 y-y1 = 3(x -x1) y-2 = 3(x+1)

OpenStudy (anonymous):

OR y = 3x +5

OpenStudy (anonymous):

ok so you subtracted ?

OpenStudy (anonymous):

yeah, and distributed the 3

OpenStudy (anonymous):

does it matter what p[oint i use to graph?